OpenStudy (anonymous):

Okay, here's the deal: I have a substance that reduced in number from 50, to 28, to 17, to 5, and then to 2. In percentages, it went from being 100% present, to 56%, then 34%, 16%, 10%, and finally 4%. What is the half-life? The substance was pieces of candy- I'm supposed to pretend it was a chemical or something going through radioactive decay. The period of time between each loss of candy was 10 seconds consistently. Could you write out your work, I want to understand how to solve this, not just get an answer. Thanks so much!!

OpenStudy (aaronq):

use the first-order integrated rate law to find the decay constant: \(\large ln[A]_t=-kt+ln[A]_o\) where \(ln[A]_o\) is the initial concentration of the substance \(k\) is the decay constant \(t\) is time \(ln[A]_t\) is the concentration of substance at time \(t\) then use: \(\large t_{1/2}=\dfrac{ln(2)}{k}\) to find the half-life after you've found \(k\). \(t_{1/2}\) is the half-life